# Performance studies of the CE-65v2 MAPS prototype structure

### v1

@@ Introduction
The Future Circular Collider (FCC [2]) is a proposed successor to the LHC. In its first phase, the FCC-ee would collide electrons and positrons at unprecedented luminosities at centre-of-mass energies between and . The requirements for FCC-ee vertex detectors are similar to the ones of ITS3 but demand further development in all areas, as shown in Table 1.
@@
For the FCC-ee vertex detector, the first layer is envisioned to be at a radius and would potentially have to deal with a much larger particle hit rate while keeping the power density comparable. A jump in resolution is also foreseen going to only sensor spatial resolution. The requirement on the material budget target in terms of % of radiation length () per layer is more relaxed, but FCC-ee vertex detectors would greatly profit from a material budget as low as ITS3 [4].
@@ 1 The CE-65v2 MAPS prototype structure
The CE-65 was developed in the context of ITS3 developments. The chip combines an analogue readout with a large pixel matrix. It is used to study the charge collection properties of the 65 nm process and compare three different amplification schemes [5]. The second version of the CE-65 prototype structure, the CE-65v2, was designed by IPHC Strasbourg and aims to go beyond the needs of ITS3. It aims to reach spatial resolution to meet the requirement of, among others, FCC-ee vertex detectors. The sensor features a large pixel matrix of pixels read out by an analogue rolling shutter readout, where the matrix is read out row-by-row. The in-pixel circuitry consists of AC-coupled amplifiers, DC-separated from the input stage of the readout electronics. The reverse bias is, thus, not limited by the electronics power supply [6]. The chip comes in 15 different variants, but in the following, the focus is on pixel pitches of 15 and , both produced in the Standard or Modified with Gap with a classic, squared, pixel arrangement.
@@ 2 Testbeam setup and analysis
The CE-65v2 chip was characterised in a test beam at the CERN SPS north area facility (H6 beamline [9]). A mixed hadron beam of was shot through the beam telescope holding the CE-65v2 device under test (DUT), six ALPIDE reference planes, and a DPTS chip for triggering22footnotetext: The authors thank the ALICE collaboration for providing the telescope and the associated DAQ software.. The CE-65v2 chips were kept at a stable temperature of using a chiller. The telescope resolution was evaluated using a telescope optimiser333https://mmager.web.cern.ch/telescope/tracking.html to achieve a resolution of .
@@ 3 Efficiency and resolution studies
The efficiency and spatial resolution study uses the cluster method to reconstruct the DUT hits. Seed thresholds ranging from to are chosen, with the lowest threshold set equal to three times the noise RMS. The achieved global efficiencies are shown in Figures 1a and 1b. All process, pitch, and reverse bias combinations yield efficiencies given sufficiently low thresholds. As expected, the STD process’s efficiency decreases rapidly with increasing seed threshold. In both processes, the sensor performs better at reverse bias compared to , indicating more depletion. For the STD process, the pitch perform significantly better than the larger pitch at high thresholds. This comes from the fact that the undepleted region at the pixel boundaries (where diffusion dominates) is wider for larger pitches [8]. This effect is not present in the GAP process, and thus, the pitch sensors feature a higher efficiency than their lower-pitch counterparts.
Figures 1c and 1d show the spatial resolutions. The STD process achieves the target resolution of at a seed threshold of , corresponding to efficiencies of for both pitches. The GAP process does not achieve such a small resolution, but () can be safely achieved at reasonable thresholds of for a pitch of (), where efficiencies are . Comparing the reverse bias voltages, it can be seen that provides better resolution in the STD process and worse resolution in the GAP process, indicating a smaller depletion depth. In the STD process, this leads to larger charge sharing and, thus, better resolution, while the GAP sensors do not benefit from this since the gap in the low-dose n-type implant inhibits charge sharing.
@@ 4 In-pixel efficiency and mean absolute deviation studies
The in-pixel measurement of the mean absolute deviation between the reconstructed telescope and DUT cluster positions () uses the window method. This way, the full benefit of the charge sharing on the resolution is considered. Figure 3 shows the mean absolute deviation depending on the position of the hit inside the pixel using the window method. In the STD process, the distribution is more uniform and does not exceed anywhere in the pixel. For the GAP sensor, the mean absolute deviation strongly deteriorates at the pixel edges and corners, which leads to the overall worse spatial resolution observed before. Using the window method, the full power of the STD process is showcased: Charge sharing from the seed to the neighbouring pixels strongly helps to reconstruct the hit position and reach resolutions much below .
@@
This work builds on previous measurements of the CE-65v2 [8] and investigated operation at reverse bias voltages of and evaluated the performance depending on the in-pixel hit position. Two different paths towards superior spatial resolution in CE-65 structures were explored. Using the STD process with large charge sharing, resolutions () are achieved with a pitch of (), while still operating with over efficiency. The GAP process, featuring faster charge collection but less charge sharing, enables resolutions of . Given a smaller pitch of 13–, it is conceivable that also the GAP process could enable sensors that fulfil the FCC-ee vertex detector spatial resolution requirement.
The challenges of achieving resolution in a large, digitally read-out chip efficient at reasonable noise levels and in a radiation environment differ between the two processes. In STD, an intricate analogue-to-digital conversion scheme is necessary to fully benefit from the charge sharing information while keeping the readout data rate minimal. For the GAP process, the challenge is to integrate all the necessary logic into pixels of small pitch required to achieve resolution.
The next steps in characterising the CE-65 are to measure the effect of irradiation and test 18 pitch sensors and the Modified without Gap process. Lastly, the impact of a staggered pixel arrangement, mimicking hexagonal pixels, is also still being evaluated.
CE-65 marks a promising step towards the DRD3 project OCTOPUS444Optimised CMOS Technology for Precision in Ultra-thin Silicon (https://octopus.web.cern.ch/), which aims to produce MAPS fulfilling the requirements of future lepton collider vertex detectors.

### v2

@@ Introduction
The Future Circular Collider (FCC [2]) is a proposed successor to the LHC. In its first phase, the FCC-ee would collide electrons and positrons at unprecedented luminosities at centre-of-mass energies between and . The requirements for FCC-ee vertex detectors (VXDs) are similar to the ones of ITS3 but demand further development in all areas, as shown in Table 1.
@@
For the FCC-ee VXD, the first layer is envisioned to be at a radius and would potentially have to deal with a much larger particle hit rate while keeping the power density comparable. A jump in resolution is also foreseen going to only sensor spatial resolution. The requirement on the material budget target in terms of % of radiation length () per layer is more relaxed, but FCC-ee VXDs would greatly profit from a material budget as low as ITS3 [4].
@@ 1 The CE-65v2 MAPS prototype structure
The CE-65 was developed in the context of ITS3 developments. The chip combines an analogue readout with a large pixel matrix. It is used to study the charge collection properties of the 65 nm process and compare three different amplification schemes [5]. The second version of the CE-65 prototype structure, the CE-65v2, was designed by IPHC Strasbourg and aims to go beyond the needs of ITS3. It aims to reach spatial resolution to meet the requirement of, among others, FCC-ee VXDs. The sensor features a large pixel matrix of pixels read out by an analogue rolling shutter readout, where the matrix is read out row-by-row. The in-pixel circuitry consists of AC-coupled amplifiers, DC-separated from the input stage of the readout electronics. The reverse bias is, thus, not limited by the electronics power supply [6]. The chip comes in 15 different variants, but in the following, the focus is on pixel pitches of 15 and , both produced in the Standard or Modified with Gap with a classic, squared, pixel arrangement.
@@ 2 Testbeam setup and analysis
The CE-65v2 chip was characterised in a test beam at the CERN SPS north area facility (H6 beamline [9]). A mixed hadron beam of was shot through the beam telescope holding the CE-65v2 device under test (DUT), six ALPIDE reference planes, and a DPTS chip also produced in [10] for triggering22footnotetext: The authors thank the ALICE collaboration for providing the telescope and the associated DAQ software.. The DUT was kept at a stable temperature of using a chiller. The telescope resolution was evaluated using a telescope optimiser333https://mmager.web.cern.ch/telescope/tracking.html to achieve a resolution of .
@@ 3 Efficiency and resolution studies
The efficiency and spatial resolution study uses the cluster method to reconstruct the DUT hits. Seed thresholds ranging from to are chosen, with the lowest threshold set equal to three times the noise RMS. The achieved global efficiencies are shown in Figures 1(a) and 1(b). All process, pitch, and reverse bias combinations yield efficiencies given sufficiently low thresholds. As expected, the STD process’s efficiency decreases rapidly with increasing seed threshold. In both processes, the sensor performs better at reverse bias compared to , indicating more depletion. For the STD process, the pitch perform significantly better than the larger pitch at high thresholds. This comes from the fact that the undepleted region at the pixel boundaries (where diffusion dominates) is wider for larger pitches [8]. This effect is not present in the GAP process, and thus, the pitch sensors feature a higher efficiency than their lower-pitch counterparts.
Figures 1(c) and 1(d) show the spatial resolutions. The STD process achieves the target resolution of at a seed threshold of , corresponding to efficiencies of for both pitches. The GAP process does not achieve such a small resolution, but () can be safely achieved at reasonable thresholds of for a pitch of (), where efficiencies are . Comparing the reverse bias voltages, it can be seen that provides better resolution in the STD process and worse resolution in the GAP process, indicating a smaller depletion depth. In the STD process, this leads to larger charge sharing and, thus, better resolution, while the GAP sensors do not benefit from this since the gap in the low-dose n-type implant inhibits charge sharing.
@@ 4 In-pixel efficiency and mean absolute deviation studies
The in-pixel measurement of the mean absolute deviation between the reconstructed telescope and DUT cluster positions () uses the window method. Figure 3 shows the mean absolute deviation depending on the position of the hit inside the pixel. In the STD process, the distribution is more uniform and does not exceed anywhere in the pixel. For the GAP sensor, the performance strongly deteriorates at the edges and corners, which leads to the overall worse spatial resolution observed before. Using the window method, the full power of the STD process is showcased: Charge sharing from the seed to the neighbouring pixels strongly helps to reconstruct the hit position and reach resolutions much below .
@@
This work builds on previous measurements of the CE-65v2 [8] and investigated operation at reverse bias voltages of and evaluated the performance depending on the in-pixel hit position. Two different paths towards superior spatial resolution in CE-65 structures were explored. Using the STD process, resolutions () are achieved with a pitch of (), while still operating with over efficiency. The GAP process, featuring faster charge collection but less charge sharing, enables resolutions of . A further reduction of the pitch could also allow the GAP process to fulfil the FCC-ee VXD spatial resolution requirement.
The challenges of achieving resolution in a l large digital read-out chip efficient at reasonable noise levels and in a radiation environment differ between the two processes. In STD, an intricate analogue-to-digital conversion scheme is necessary to fully benefit from the charge sharing information while keeping the readout data rate minimal. For the GAP process, the challenge is to integrate all the necessary logic into pixels of small pitch required to achieve resolution.
The next steps in characterising the CE-65 are to measure the effect of irradiation and test pitch sensors and the Modified without Gap process. Lastly, the impact of a staggered pixel arrangement, mimicking hexagonal pixels, is also still being evaluated.
CE-65 marks a promising step towards MAPS fit for future lepton collider vertex detectors.